OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? s + (-10) = 14

OpenStudy (anonymous):

s - 10 = 14 s = 14 + 10 s = 24

OpenStudy (zzr0ck3r):

s+(-10) = 14 if we add 10 to both sides we get s+(-10) +10 = 14 + 10 then simplify s-10+10 = 24 s= 24 so 10 is the number to add to both sides to solve for s

OpenStudy (anonymous):

lol there we go 3 answers :)